# Unsupervised Log Anomaly Detection with Few Unique Tokens

### v3

@@ Abstract
[…]
1 Lawrence Berkeley National Lab (LBNL), 1 Cyclotron Rd, Berkeley, CA 94720, USA2 Deutsches Elektronen-Synchrotron (DESY), Notkestr. 85, 22607 Hamburg, Germany 3 Technische Universitaet Hamburg (TUHH), Am Schwarzenberg-Campus 1, 21073, Hamburg, Germany
@@ 2 Related Work
A common initial approach to log anomaly detection involves manually defined rule-based systems. For instance, Cinque et al. [3] and Yen et al. [20] developed methods that scan logs for predefined patterns indicative of anomalies. However, these approaches rely heavily on expert knowledge to construct and maintain effective rules, a labor-intensive process that struggles with evolving systems. This limitation has spurred the development of more automated techniques leveraging machine learning.
With the rise of machine learning (ML), deep learning-based approaches have shown potential for thorough log analysis, particularly when a large log corpus is available, often accompanied by extensive, manually created labels. Long Short-Term Memory (LSTM) recurrent neural networks [6, 21, 22] have been popular for log anomaly detection due to their ability to handle sequential data. More recently, Transformer architectures [18], such as BERT [5], have also been applied for this task [9]. While powerful, the reliance of these deep learning models on vast training datasets and millions of parameters can limit their applicability in resource-constrained scenarios like ours at EuXFEL, where such large, labeled datasets are not readily available. For a broader survey of ML in log analysis, refer to [10].
Addressing the need for automated feature representation, Bertero et al. [1] proposed treating logs as natural text, leveraging Word2Vec word vector representations [14, 13] for automated embedding. This technique maps words to a vector space, allowing the use of off-the-shelf classifiers for anomaly detection. While this automates embedding, a major drawback is that their approach still relies on manual labeling to train the classifier, which can be prohibitively expensive. Furthermore, they treat each log entry independently, thereby ignoring the crucial sequential relationships between consecutive log messages.
To mitigate the need for labeled data, other works, such as [11] and [19], have explored unsupervised learning. These methods typically apply text mining techniques to logs and employ clustering algorithms to identify anomalies without manual labels. However, a common limitation is that they, too, often consider log entries in isolation rather than fully leveraging the contextual information embedded in log sequences.
@@
Inspired by the embedding techniques in [1], we employ word vectors to represent log entries in a high-dimensional space, which helps mitigate data scarcity. However, crucially, instead of relying on supervised classifiers, we model the logs from each source as a sequence of events. Our key insight is to focus on modeling the patterns of co-occurring log sequences using an unsupervised technique based on Hidden Markov Models (HMMs) operating on these embedded sequences. By learning these sequential regularities, anomalies are detected as deviations from established contextual patterns, rather than solely from individual log content. This probabilistic HMM-based approach requires estimating only a minimal number of parameters, enabling robust anomaly detection even with limited and unlabeled training data.
@@ 3 Method
Normal Last event has unusual position Anomalous event
Refer to caption Refer to caption Refer to caption
Figure 1: HMM Anomaly Detection Performance on a Synthetic Log Sequence: This figure demonstrates the proposed HMM’s ability to detect anomalies in a synthetic log sequence. The sequence involves normal events ($o_{1}$,$o_{2}$) and a predefined anomalous event ($o_{a}$). For each scenario, the HMM parameters were estimated using the sequence except the evaluated entry (to ensure the model scores novel or altered final states without prior knowledge of them, minimizing overfitting for this test). The resulting anomaly score ($s$) for the final entry is observed: Left Panel (Normal Pattern): A sequence ending with the established, repetitive $o1$, $o2$ pattern. Value of $s$ remains consistently low, correctly identifying the sequence as normal. Center Panel (Disrupted Pattern): The sequence concludes with a swapped order ($o_{2}$, $o_{1}$), disrupting the expected $o_{1}$, $o_{2}$ pattern. This unexpected sequencing results in a noticeable increase in the $s$. Right Panel (Anomalous Event): The sequence terminates with the insertion of the novel, anomalous event $o_{a}$. This causes a substantial spike in the $s$, clearly highlighting the improbable observation. These scenarios illustrate that the HMM effectively detects anomalies stemming from both unexpected sequences of otherwise normal events (center) and the introduction of entirely novel or improbable log messages (right). Increases in the $s$ directly correlate with the degree of deviation from learned sequential patterns.
@@ 3.1 Preprocessing and Tokenization
  • Tokenization: Each segmented log entry is then tokenized into its constituent words or symbols using the NLTK punkt tokenizer [2].
@@
      Numeric Abstraction: Numeric tokens are categorized and replaced with placeholders to generalize numerical values: non-zero numbers become $nzandzerosarereplacedwith$zero.
@@
      Stopword Removal: Common English stopwords (e.g., ”is”, ”a”, ”the”), as defined by the NLTK library [2], are removed to reduce noise and focus on more content-rich tokens. Embedding
@@ 3.2 Embedding
To represent log entries as dense vectors, our approach utilizes Word2Vec [14, 13]. The fundamental principle of Word2Vec is that words appearing in similar textual contexts are likely to possess similar semantic meanings. It operationalizes this by training a shallow neural network to learn word embeddings that capture these contextual semantic relationships. For this purpose, we employ the Continuous Bag-of-Words (CBOW) architecture, as introduced in [13], which learns to predict a target word from its surrounding context words (an alternative to the skip-gram model, which predicts the context from a target word).
@@
This additive property is particularly valuable for generating a single vector representation for multi-word log entries. In our method, we achieve this by calculating the mean of the Word2Vec embeddings for all tokens within a given log entry. While more complex pooling strategies exist [7, 16], this mean pooling approach proved both effective and sufficient for our requirements, providing a concise yet semantically rich representation for each log message.
Refer to caption Refer to caption Refer to caption
Figure 2: (Left) A 2D UMAP [12] projection of the 16-dimensional Word2Vec embeddings for all words in the vocabulary. Semantically similar words form distinct clusters, highlighted by red ellipses (or markers/color, specify how they are distinguished). Examples include clusters for three-letter month names, and operating system terms (e.g., ’windows’, ’linux’, ’mac’, ’os’). This demonstrates that the Word2Vec model successfully captured latent semantic relationships even under these very limited conditions. (Center) A 2D UMAP projection of unique log entries, where each entry is represented by the sum of its 16-dimensional token embeddings. The non-uniform distribution indicates that log messages predominantly form a few distinct clusters, rather than being widely dispersed. (Right) Frequency distribution of unique, pre-processed log entries. The distribution exhibits a strong positive skew, indicating that a very limited set of messages accounts for the vast majority of log occurrences.
@@ 3.3 Anomaly detection with HMMs
Adopting notation similar to [4], let $Q=\{q_{1},...,q_{N}\}$ represent a set of $N$ hidden states, and $O=\left(o_{1},...,o_{T}\right)$ denote a sequence of $T$ observations. HMMs are characterized by two fundamental assumptions:
@@
A key requirement in log anomaly detection is the ability to handle novel entries [22]. While our method’s generalization to logs containing entirely new, out-of-vocabulary terms is constrained by the initial embedding vocabulary, its emphasis on sequence modeling over individual word semantics is a core strength. This focus allows it to detect anomalies primarily through contextual irregularities and unexpected variations in log sequences, rather than relying solely on the content of individual messages. Indeed, we have observed that anomalies in our domain often manifest as unusual sequences of known events rather than just the appearance of specific unknown terms. By basing the anomaly score si on the likelihood of an observed sequence under the learned HMM (where significant deviations from expected sequences, implying lower likelihoods, result in higher anomaly scores), rather than on predefined keyword rules, our approach can effectively evaluate previously unseen sequences of known log messages for contextual anomalies. This unsupervised, sequential methodology, which assesses log entries based on their contextual fit within learned patterns rather than relying on predefined labels, distinguishes our technique from supervised classification approaches such as those in [1, 9, 6, 21, 22]. However, it is important to acknowledge that false alarms may still occur if natural system fluctuations also produce sequences that deviate significantly from the learned patterns, an aspect we illustrate further in 7
@@ 3.5 Parameter Estimation
For the real-world results presented (Figures 456, and 7), HMM parameters were consistently estimated using all available log messages that occurred prior to a designated test period. The log messages within this subsequent period were then exclusively used as the test set, ensuring a strict separation between training and evaluation data.
@@ 3.6 Handling Long and Noisy Sequences
Normal Last event has unusual position Anomalous event
Refer to caption Refer to caption Refer to caption
@@ 3.9 Implementation
The implementation was developed using Python version 3.9. For the generation of log message embeddings via Word2Vec, we utilized the Gensim library [15]. Sequence modeling with Hidden Markov Models was performed using the Hmmlearn package[17].
@@ 4 Experiments and Results
To demonstrate the performance of our anomaly detection method, this section presents an analysis of four representative instances. The input logs and their corresponding computed anomaly scores for these instances are illustrated in Figures 456, and 7.
@@
Score of 18 events Score of 100 events
Refer to caption Refer to caption
@@
Score of 20 events Score of 100 events
Refer to caption Refer to caption
@@
Score of 26 events Score of 100 events
Refer to caption Refer to caption
@@
Score of 26 events Score of 100 events
Refer to caption Refer to caption
@@ 5 Future Work
One immediate direction is the exploration of more advanced or complementary anomaly detection techniques, such as those by [8], which could potentially offer improved precision or recall. Furthermore, enhancing the verbosity of log data, if feasible, could generate richer datasets. This, in turn, might enable the effective deployment of more complex, parameter-rich algorithms, addressing a common challenge with sparse log information highlighted in related work.

### v4

@@ 2 Related Work
A common initial approach to log anomaly detection involves manually defined rule-based systems. For instance, Cinque et al. Cinque et al., (2012) and Yen et al. Yen et al., (2013) developed methods that scan logs for predefined patterns indicative of anomalies. However, these approaches rely heavily on expert knowledge to construct and maintain effective rules, a labor-intensive process that struggles with evolving systems. This limitation has spurred the development of more automated techniques leveraging machine learning.
With the rise of machine learning (ML), deep learning-based approaches have shown potential for thorough log analysis, particularly when a large log corpus is available, often accompanied by extensive, manually created labels. Long Short-Term Memory (LSTM) recurrent neural networks Du et al., (2017); Zhang et al., (2016, 2019) have been popular for log anomaly detection due to their ability to handle sequential data. More recently, Transformer architectures Vaswani et al., (2017), such as BERT Devlin et al., (2018), have also been applied for this task Guo et al., (2021). While powerful, the reliance of these deep learning models on vast training datasets and millions of parameters can limit their applicability in resource-constrained scenarios like ours at EuXFEL, where such large, labeled datasets are not readily available. For a broader survey of ML in log analysis, refer to Landauer et al., (2023).
Addressing the need for automated feature representation, Bertero et al. Bertero et al., (2017) proposed treating logs as natural text, leveraging Word2Vec word vector representations Mikolov et al., 2013b ; Mikolov et al., 2013a for automated embedding. This technique maps words to a vector space, allowing the use of off-the-shelf classifiers for anomaly detection. While this automates embedding, a major drawback is that their approach still relies on manual labeling to train the classifier, which can be prohibitively expensive. Furthermore, they treat each log entry independently, thereby ignoring the crucial sequential relationships between consecutive log messages.
To mitigate the need for labeled data, other works, such as Lou et al., (2010) and Xu et al., (2009), have explored unsupervised learning. These methods typically apply text mining techniques to logs and employ clustering algorithms to identify anomalies without manual labels. However, a common limitation is that they, too, often consider log entries in isolation rather than fully leveraging the contextual information embedded in log sequences.
@@
Inspired by the embedding techniques in Bertero et al., (2017), we employ word vectors to represent log entries in a high-dimensional space, which helps mitigate data scarcity. However, crucially, instead of relying on supervised classifiers, we model the logs from each source as a sequence of events. Our key insight is to focus on modeling the patterns of co-occurring log sequences using an unsupervised technique based on Hidden Markov Models (HMMs) operating on these embedded sequences. By learning these sequential regularities, anomalies are detected as deviations from established contextual patterns, rather than solely from individual log content. This probabilistic HMM-based approach requires estimating only a minimal number of parameters, enabling robust anomaly detection even with limited and unlabeled training data.
@@ 3 Method
Figure 1: HMM Anomaly Detection Performance on a Synthetic Log Sequence: This figure demonstrates the proposed HMM’s ability to detect anomalies in a synthetic log sequence. The sequence involves normal events ($o_{1}$$o_{2}$) and a predefined anomalous event ($o_{a}$). For each scenario, the HMM parameters were estimated using the sequence except the evaluated entry (to ensure the model scores novel or altered final states without prior knowledge of them, minimizing overfitting for this test). The resulting anomaly score ($s$) for the final entry is observed: Left Panel (Normal Pattern): A sequence ending with the established, repetitive $o1$, $o2$ pattern. Value of $s$ remains consistently low, correctly identifying the sequence as normal. Center Panel (Disrupted Pattern): The sequence concludes with a swapped order ($o_{2}$, $o_{1}$), disrupting the expected $o_{1}$, $o_{2}$ pattern. This unexpected sequencing results in a noticeable increase in the $s$. Right Panel (Anomalous Event): The sequence terminates with the insertion of the novel, anomalous event $o_{a}$. This causes a substantial spike in the $s$, clearly highlighting the improbable observation. These scenarios illustrate that the HMM effectively detects anomalies stemming from both unexpected sequences of otherwise normal events (center) and the introduction of entirely novel or improbable log messages (right). Increases in the $s$ directly correlate with the degree of deviation from learned sequential patterns.
@@ 3.1 Preprocessing and Tokenization
  • Tokenization: Each segmented log entry is then tokenized into its constituent words or symbols using the NLTK punkt tokenizer Bird et al., (2009).
@@
      Numeric Abstraction: Numeric tokens are categorized and replaced with placeholders to generalize numerical values: non-zero numbers become \$nz and zeros are replaced with \$zero.
@@
      Stopword Removal: Common English stopwords (e.g., ”is”, ”a”, ”the”), as defined by the NLTK library Bird et al., (2009), are removed to reduce noise and focus on more content-rich tokens. Embedding
@@ 3.2 Embedding
To represent log entries as dense vectors, our approach utilizes Word2Vec Mikolov et al., 2013b ; Mikolov et al., 2013a . The fundamental principle of Word2Vec is that words appearing in similar textual contexts are likely to possess similar semantic meanings. It operationalizes this by training a shallow neural network to learn word embeddings that capture these contextual semantic relationships. For this purpose, we employ the Continuous Bag-of-Words (CBOW) architecture, as introduced in Mikolov et al., 2013a , which learns to predict a target word from its surrounding context words (an alternative to the skip-gram model, which predicts the context from a target word).
@@
This additive property is particularly valuable for generating a single vector representation for multi-word log entries. In our method, we achieve this by calculating the mean of the Word2Vec embeddings for all tokens within a given log entry. While more complex pooling strategies exist Gao et al., (2021); Reimers and Gurevych, (2019), this mean pooling approach proved both effective and sufficient for our requirements, providing a concise yet semantically rich representation for each log message.
Figure 2: (Left) A 2D UMAP McInnes et al., (2018) projection of the 16-dimensional Word2Vec embeddings for all words in the vocabulary. Semantically similar words form distinct clusters, highlighted by red ellipses (or markers/color, specify how they are distinguished). Examples include clusters for three-letter month names, and operating system terms (e.g., ’windows’, ’linux’, ’mac’, ’os’). This demonstrates that the Word2Vec model successfully captured latent semantic relationships even under these very limited conditions. (Center) A 2D UMAP projection of unique log entries, where each entry is represented by the sum of its 16-dimensional token embeddings. The non-uniform distribution indicates that log messages predominantly form a few distinct clusters, rather than being widely dispersed. (Right) Frequency distribution of unique, pre-processed log entries. The distribution exhibits a strong positive skew, indicating that a very limited set of messages accounts for the vast majority of log occurrences.
@@ 3.3 Anomaly detection with HMMs
Adopting notation similar to Daniel et al., (2007), let $Q=\{q_{1},...,q_{N}\}$ represent a set of $N$ hidden states, and $O=\left(o_{1},...,o_{T}\right)$ denote a sequence of $T$ observations. HMMs are characterized by two fundamental assumptions:
@@
A key requirement in log anomaly detection is the ability to handle novel entries Zhang et al., (2019). While our method’s generalization to logs containing entirely new, out-of-vocabulary terms is constrained by the initial embedding vocabulary, its emphasis on sequence modeling over individual word semantics is a core strength. This focus allows it to detect anomalies primarily through contextual irregularities and unexpected variations in log sequences, rather than relying solely on the content of individual messages. Indeed, we have observed that anomalies in our domain often manifest as unusual sequences of known events rather than just the appearance of specific unknown terms. By basing the anomaly score si on the likelihood of an observed sequence under the learned HMM (where significant deviations from expected sequences, implying lower likelihoods, result in higher anomaly scores), rather than on predefined keyword rules, our approach can effectively evaluate previously unseen sequences of known log messages for contextual anomalies. This unsupervised, sequential methodology, which assesses log entries based on their contextual fit within learned patterns rather than relying on predefined labels, distinguishes our technique from supervised classification approaches such as those in Bertero et al., (2017); Guo et al., (2021); Du et al., (2017); Zhang et al., (2016, 2019). However, it is important to acknowledge that false alarms may still occur if natural system fluctuations also produce sequences that deviate significantly from the learned patterns, an aspect we illustrate further in 7
@@ 3.5 Parameter Estimation
For the real-world results presented (Figures 4, 5, 6, and 7), HMM parameters were consistently estimated using all available log messages that occurred prior to a designated test period. The log messages within this subsequent period were then exclusively used as the test set, ensuring a strict separation between training and evaluation data.
@@ 3.9 Implementation
The implementation was developed using Python version 3.9. For the generation of log message embeddings via Word2Vec, we utilized the Gensim library Řehůřek and Sojka, (2010). Sequence modeling with Hidden Markov Models was performed using the Hmmlearn packageSergei Lebedev, (2023).
@@ 4 Experiments and Results
To our knowledge, there is no standardized dataset currently available for this specific purpose. Due to privacy constraints, we are unable to release our full datasets, with the exception of a few minimal examples. All examples included in this paper have been thoroughly vetted to ensure that no private or sensitive information is revealed. To showcase the efficacy of our proposed method, we have carefully chosen and presented a number of representative examples. The evaluation of these examples was conducted by human operators.
For evaluation, we used a setup with Intel(R) Core(TM) i7-11850H @ 2.50GHz and 16GB. No acceleration was used.
To demonstrate the performance of our anomaly detection method, this section presents an analysis of four representative instances. The input logs and their corresponding computed anomaly scores for these instances are illustrated in Figures 4, 5, 6, and 7.
@@ 5 Future Work
One immediate direction is the exploration of more advanced or complementary anomaly detection techniques, such as those by Görnitz et al., (2015), which could potentially offer improved precision or recall. Furthermore, enhancing the verbosity of log data, if feasible, could generate richer datasets. This, in turn, might enable the effective deployment of more complex, parameter-rich algorithms, addressing a common challenge with sparse log information highlighted in related work.
@@ 6 Conclusion
The code and materials for this study are openly available in a GitHub repository at https://github.com/sulcantonin/LOG_ICALEPCS23.
@@ 7 Acknowledgement
An Grammarly and Gemini AI-powered language tool was employed to enhance the manuscript’s grammar and clarity, without being used for text generation.